OpenStudy (anonymous):

Evaluate 9x2y-2 for x = –3 and y = 2.

OpenStudy (anonymous):

for x=-3,y=2, 9*-3*2*2-2=-108-2=-110

OpenStudy (anonymous):

is it clear?

OpenStudy (anonymous):

that's not even a answer Choice

OpenStudy (anonymous):

is it not 9*x*2*y-2?

OpenStudy (anonymous):

http://oi50.tinypic.com/33ndc74.jpg

OpenStudy (hba):

@KinqMiquel Did you try to plug in the values for x and y and solve it for yourself ? If Yes,What did you get ?

OpenStudy (anonymous):

No i didn't because i Don't Know Mathematics..

OpenStudy (hba):

@Praja_01 Giving direct answer is not allowed on Os,Please read the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok well thanks for your time bye

OpenStudy (anonymous):

srry hba,

OpenStudy (hba):

@KinqMiquel Try to solve the question on your own and we will tell you where you go wrong.

OpenStudy (hba):

@Praja_01 No problem,Thanks a lot to understand.